KORBIN COMPANY
Comparative Balance Sheets
December 31, 2019, 2018, and 2017
2019 2018 2017
Assets
Current assets $ 52,390 $ 37,924 $ 51,748
Long-term investments 0 500 3,950
Plant assets, net 100,000 96,000 60,000
Total assets $ 152,390 $ 134,424 $ 115,698
Liabilities and Equity
Current liabilities $ 22,800 $ 19,960 $ 20,300
Common stock 72,000 72,000 60,000
Other paid-in capital 9,000 9,000 6,000
Retained earnings 48,590 33,464 29,398
Total liabilities and equity $ 152,390 $ 134,424 $ 115,698

3. Complete the below table to calculate the balance sheet data in trend percents with 2017 as base year. (Round your percentage answers to 2 decimal places.)

KORBIN COMPANY
Balance Sheet Data in Trend Percents
December 31, 2019, 2018 and 2017
2019 2018 2017
Assets
Current assets % % 100.00 %
Long-term investments 100.00
Plant assets, net 100.00
Total assets % % 100.00 %
Liabilities and Equity
Current liabilities % % 100.00 %
Common stock 100.00
Other paid-in capital 100.00
Retained earnings 100.00
Total liabilities and equity % % 100.00 %
Answer #1

The balance sheet data in trend percents with base year 2017 is calculated as:

2017 % Korbin Company Balance sheet Data in Trend Percents December 31, 2019,2018,2017 2019 % 2018 % Assets Current assets 10

Working notes:

J K 2017% F G H I Korbin Company Balance sheet Data in Trend Percents December 31,2019,2018,2017 2019% 2018 % 8 Assets 9 Curr

2019 2017 % Korbin Company Balance sheet Data in Trend Percents December 31,2019,2018,2017 % 2018 Assets Current assets =F9/1

2017% Korbin Company Balance sheet Data in Trend Percents December 31,2019,2018,2017 2019 % 2018 % Assets Current assets 101.24 % 72.07% Long-term investments 0 % 12.6582278 % Plant assets, net 166.67% 160.00% Total assets 131.71 % 116.19 % Liabilities and Equity Current liabilities 112.32 % 98.33% Common stock 120.00% 120% Other paid in capital 150.00% 150.00% Retained earnings 165.28 % 113.83 % Total liabilities 131.71 % 116.19% 100% 100% 100% 100% 100% 100% 100% 100% 100%

J K 2017% F G H I Korbin Company Balance sheet Data in Trend Percents December 31,2019,2018,2017 2019% 2018 % 8 Assets 9 Current assets 52,390 37,294 10 Long-term investments 500 11 Plant assets, net 100,000 96,000 12 Total assets 152,390 134,424 13 Liabilities and Equity 14 Current liabilities 22,800 19,960 15 Common stock 72,000 72,000 16 Other paid in capital 9,000 17 Retained earnings 48,590 33,464 18 Total liabilities 152,390 134,424 51,748 3,950 60,000 115,698 9,000 20,300 60,000 6,000 29,398 115,698

2019 2017 % Korbin Company Balance sheet Data in Trend Percents December 31,2019,2018,2017 % 2018 Assets Current assets =F9/19*100 =H9/19*100 Long-term investments =F10/110*100 =H10/110*100 Plant assets,net =F11/111*100 =H11/J11*100 Total assets =F12/112*100 =H12/112*100 Liabilities and Equity Current liabilities =F14/114*100 =H14/114*100 Common stock =F15/315*100 =H15/115*100 Other paid in capital =F16/116*100 % =H16/116*100 Retained earnings =F17/117*100 =H17/117*100 Total liabilities =F18/H18*100 =H18/118*100 %